OpenStudy (anonymous):

Given the constraints, find the vertices of the feasible region and the maximum value of the function: f(x,y)=x-y x + 2y < 6 2x - y < 7 x > -2 y > -3 What is the maximum?

OpenStudy (freckles):

we need to figure out the vertices from the inequalities there

OpenStudy (freckles):

so i would graph the inequalities first

OpenStudy (anonymous):

i got it now . can u help me with another question tho

OpenStudy (anonymous):

Which ordered triple is a solution of the following linear system?: 3n = -18 3m - 2n + 4p = 12 2m - 4n = 16

OpenStudy (freckles):

ok you can find n first from the first equation once you find n from that equation you find m from the equation 2m-4n=16

OpenStudy (freckles):

so 3n=-18 what is n?